OpenStudy (anonymous):

What is the length of the diagonal of a square with a 15-decimeter side?

OpenStudy (anonymous):

so if you make a diagonal that would give us two triangles right? c^2=a^2+b^2 c^2=225+225 sqr root= 21.21 decimeter

OpenStudy (anonymous):